OpenStudy (anonymous):

:( CAN ANYONE PLEASE HELP!!!!!!!!!!!!!!!!!!!!! For questions 3 and 4, order the group of quadratic functions from widest to narrowest graph. 3.) y=3x^2, y=x^2, y=2x^2 4.) y=-5x^2, y=-x6^2, y=-4x^2

OpenStudy (anonymous):

@Tears_of_Ashes can u help:(

OpenStudy (anonymous):

I can help

OpenStudy (anonymous):

what am i finding exactly?

OpenStudy (anonymous):

well its multipule choice for both but i think they want it in order idk:(

OpenStudy (anonymous):

whats the multiple choice ?

OpenStudy (anonymous):

give me a min k

OpenStudy (anonymous):

I can still help

OpenStudy (anonymous):

3. A. y=2x^2, y=x^2, y=3x^2 B. y=3x^2, y=2x^2,y=x^2 C. y=x^2,y=2x^2,y=3x^2 D. y=x^2, y=3x^2,y=2x^2

OpenStudy (anonymous):

sure @Darius25 whoever can help

OpenStudy (anonymous):

4. A. y=-x^2,y=-5x^2,y=-4x^2 B. y=-x^2,y=-4x^2,y=-5x^2 C. y=-5x^2,y=-4x^2,y=-x^2 D. y=-4x6^2,y=-x^2,y=-5x^2

OpenStudy (anonymous):

Here @Tears_of_Ashes

OpenStudy (anonymous):

whats the 4 for? is that what its suppose to equal?

OpenStudy (anonymous):

sorry im just trying to understand the question

OpenStudy (anonymous):

theres #3 and #4

OpenStudy (anonymous):

okay im working on it

OpenStudy (anonymous):

what grade are you in? cuz im not sure i know this..

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

Its ok i cann guess but im scaared because i dont undertannnndddd either:(

OpenStudy (anonymous):

I can do it I am advanced 7th grade it is very simple

OpenStudy (anonymous):

i would say go with C y=x^2, y=2x^2, y=3x^2 but thats if you have to put it in order but i dont want you to be wrong... :(

OpenStudy (anonymous):

and @Darius25 if you know how to do it help the girl dam

OpenStudy (anonymous):

well help me please @Darius25 ?

OpenStudy (anonymous):

thank you @Tears_of_Ashes i appreciate it!:)your awesome

OpenStudy (anonymous):

Thanks (: you are to but really if someone is going to keep saying they can help they can HELP by telliing us what the fuk to do lol @Darius25

OpenStudy (anonymous):

idk i dont think that guy will help me anywhooo lol:)

OpenStudy (anonymous):

lol really nuuh! idgaf

OpenStudy (anonymous):

ya probably not but i tried as much as i could

OpenStudy (anonymous):

just go for C for #3 and if you get it wrong you can blame it one me -.-

OpenStudy (anonymous):

OKAYN SOUNDS LIKE A DEALL:)

OpenStudy (anonymous):

YAY! And if you get it right IM AWESOME! lol :)

OpenStudy (anonymous):

How is the graph of y=4x^2 +1 different from the graph of y=4x^2? A. it shifted up 1 unit b. it shifted down 1 unit c. it shifted left 1 unit d. it shifted 1 unit right MY LAST 1 AND ILL TURN IT IN CAN U TELL ME THIS ONNEEEE PLEASE LOL:)

OpenStudy (anonymous):

YOUR AWESOME EITHER WAY @Tears_of_Ashes

OpenStudy (anonymous):

yepperz hold on a min

OpenStudy (anonymous):

OH TAYYY

OpenStudy (anonymous):

i would say it shifts up 1 but this one im really really not sure about

OpenStudy (anonymous):

i got 4 outa 5 thank u:)

OpenStudy (anonymous):

4.. was B.:)

OpenStudy (anonymous):

oh so i got them right!:D

OpenStudy (anonymous):

SO U R Awesome

OpenStudy (anonymous):

YESSSS <3